Find the integral surface of the equation 4yzp+q+2y=0 passing through

y2+z2=2,x+z=1


1
Expert's answer
2021-11-01T17:43:16-0400

Characteristic system

"\frac{dx}{4yz}=\frac{dy}{1}=\frac{dz}{-2y}"

First and last terms can be rewritted as

"\frac{dx}{4yz}=\frac{-2zdz}{4yz}"

Therefore

dx=--2zdz

"d(x+z^2)=0;
x+z^2=C_1-first\space integral"

Second and third terms of the characteristic system give:

-2ydy=dz;

"d(z+y^2)=0;
z+y^2=C_2-second\space integral"

General solution may be written in the form^

"x+z^2=F(z+y^2)"

where F(t)-some function

Let "y^2+z^2=2,z+x=1-" initial curve

Then "z+y^2=2+z-z^2,x+z^2=1-z+z^2"

Must be:

"1-z+z^2=F(2+z-z^2)=F(3-(1-z+z^2))\space \forall z"

or

F(3-t)=t=-(3-t)+3? therefore F(u)=3-u

Therefore surface equation is:

"x+z^2=3-z-y^2;\\
Answer:\\
z^2+y^2+z+x=3"